Tell whether if the following piecewise function is a continuous at a given point or not. (SHOW THE SOLUTION).



2. at x = 0



1/x if x ≤ -1


x² - 2 if x > -1




1
Expert's answer
2022-06-09T06:14:17-0400

"f(x)" is continuous at a point "x=a" if "f(a)" is defined and "lim_{x\to a} f(x)=f(a)."

The function

"f(x)=\{\begin{matrix}
 \frac{1}{x}, & \text{if }x\leq-1, \\
 x^2-2, & \text{if }x>-1,
\end{matrix}"

is defined at "x=0:"

"f(0)=0^2-2=-2."


"lim_{x\to 0^-} f(x)=lim_{x\to 0^-} (x^2-2)=0^2-2=-2,"


"lim_{x\to 0^+} f(x)=lim_{x\to 0^+} (x^2-2)=0^2-2=-2=lim_{x\to 0^-} f(x),"


so there exist "lim_{x\to 0} f(x)=-2=f(0)" and the function is continuous at "x=0."


We can see the continuity at "x=0" at the graph of "f(x)" :




Answer: the function is continuous at x=0.